# Artificial Precision Polarization Array: Sensitivity for the axion-like dark matter with clock satellites

### v2

@@ Abstract
The approaches to searching for axion-like signals based on pulsars include observations with pulsar timing arrays (PTAs) and pulsar polarization arrays (PPAs). However, these methods are limited by observational uncertainties arising from multiple unknown and periodic physical effects, which substantially complicate subsequent data analysis. To mitigate these issues and improve data fidelity, we propose the Artificial Pulsar Polarization Arrays (APPA): a satellite network comprising multiple pulsed signal transmitters and a dedicated receiver satellite. To constrain the axion-photon coupling parameter $g_{a\gamma}$, we generate simulated observations using Monte Carlo methods and investigate the sensitivity of APPA using two complementary approaches: Likelihood analysis and frequentist analysis. Simulations indicate that for the axion mass range of $10^{-22}-10^{-18}$ eV, APPA yields a tighter upper limit on $g_{a\gamma}$ (at the 95% confidence level) than conventional ground-based observations, while also achieving superior detection sensitivity. Moreover, a larger spatial distribution scale of the satellite network corresponds to a greater advantage in detecting axions with lighter masses.
@@ II Axion induced birefringence effect
In [32], the authors considered the occupation number of a scalar axion field and estimated it under standard local halo assumptions. Assuming a characteristic dark-matter velocity of $10^{-3}c$, and a local dark matter energy density $\rho_{\text{DM}}=0.4$GeV/cm3 in the vicinity of the Solar System, the occupation number can be estimated by
@@
Meanwhile, noting that the typical dark matter velocity is $v\sim 10^{-3}c$. In realistic situations, the coherence time of the axion field can be expressed as $\tau_{c}=\frac{1}{m_{a}v^{2}}$. If the condition $\tau_{c}\gg T_{\text{obs}}$ is satisfied for a given axion mass, then the observed axion field amplitude behaves stochastically. As shown in [20, 13], within a single coherence domain the amplitude distribution of the axion field follows a Rayleigh distribution,
$a_{0}=\frac{\sqrt{2\rho_{\text{DM}}}}{m_{a}}\alpha,\qquad p\big(\alpha\big)\sim\alpha e^{-\frac{\alpha^{2}}{2}}.$ (II.8)
By combining Eq. (II.4), Eq. (II.7) and Eq. (II.8), we obtains
$\Delta\phi=\phi_{a}\mathrm{cos}\big(m_{a}t+\varphi_{a}\big),$ (II.9)
@@
$\phi_{a}$ $=\frac{g_{a\gamma}}{\sqrt{2}m_{a}}\left(\rho_{o}\alpha_{o}^{2}+\rho_{s}\alpha_{s}^{2}-2\sqrt{\rho_{o}\rho_{s}}\alpha_{o}\alpha_{s}\mathrm{cos}\theta_{T}\right)^{\frac{1}{2}},$ (II.10)
with $\theta_{T}=m_{a}T_{i}+\varphi_{o}-\varphi_{s}$, $\rho_{s}$ and $\rho_{o}$ denote the dark matter energy density in the vicinity of the source and the Solar System, respectively. $T_{i}\equiv L_{i}/c$ is the signal travel time from a transmitting satellite $S_{i}$ to the receiving satellite $O$.
@@ III Methods of analysis
2. Uncertainty in local dark matter energy density. As indicated in Eq. (II.10), the measurement of the axion–photon coupling parameter $g_{a\gamma}$ requires knowledge of the dark matter energy densities $\rho_{s}$ near the pulsar and $\rho_{o}$ near the Solar System. However, $\rho_{s}$ is difficult to determine observationally and is subject to large uncertainties, which ultimately limit the precision of the derived constraints.
3. Randomness of pulsar signal phase. The pulse phase $\varphi_{s}$ of a pulsar signal is inherently random and indeterminate, which prevents the signal propagation time $T$ from being used to reliably determine the value of $\Theta$ in Eq. (II.10).
@@
To address these observational challenges at their source, we extend the concept proposed in [2] by introducing the APPA, a network of space-based satellites equipped with ultra precise clocks and capable of regularly transmitting pulsed signals to a dedicated receiving satellite. The transmitted signals are strictly periodic, with a timing uncertainty of $\sigma_{t}=\xi_{\delta t}\delta t$ [2], and carry preconfigured polarization information. As illustrated in Fig. 1, the receiving satellite, positioned at the center of the transmitting network, continuously collects APPA signals and can accurately measure both the distance to and the direction of each transmitting satellite.
@@
$\lambda=\big(EF\sigma_{\text{err}}\big)^{2}+EQ^{2}.$ (III.1)
Here, $\sigma_{\text{err}}$ denotes the instrumental measurement error, which we take to be $\sigma_{\text{err}}=0.3^{\circ}$. The $EF$ and $EQ$ represent the white noise scaling factor and the additive white noise, respectively, and their prior distributions are specified as follows:
$P\big(EF\big)$ $=\mathcal{N}\big[1,1\big],$ (III.2)
@@
First, we show the main conclusions of this work in advance. We present a comparison between the simulation results of this work and existing observational constraints in a single figure (see Fig. 2). Owing to the long observational baseline $T_{\text{obs}}$ and the highly stable signal emission cadence $\Delta t$, APPA is theoretically capable of probing an exceptionally broad range of axion induced oscillation frequencies. However, as shown in Fig. 2, although APPA theoretically possesses an extremely wide detection range, the domain where it truly demonstrates an advantage in axion signal detection is limited to $m_{a}\sim\mathcal{O}\big(10^{-22}–10^{-18}\big)$eV, both the 95% confidence upper limit $g_{95\%}$ derived from likelihood analysis and the detection sensitivity $g_{a\gamma}$ obtained via frequentist analysis highlight the significant advantages of the APPA framework over current ground-based observations.
Figure 2: Comparison between the APPA simulation results and existing observational constraints. Color coding: red, CAST [5]; magenta, PPA [57]; green, PPTA [12]; blue, likelihood analysis (L=5L=5AU), we shows that APPA has a detection advantage within $m_{a}\sim\mathcal{O}\big(10^{-22}–10^{-18}\big)$eV and exhibits comparable sensitivity to axions with different masses; purple, frequentist analysis, the detection sensitivity is set to $\phi_{a}=0.01^{\circ}$ in Eq. (III.9), consistent with current observational limits on linear polarization angle rotation, with minor fluctuations attributed to instrumental noise. The black dashed line marks the critical boundary, to the left of this boundary, the condition $\mathrm{cos}\big(m_{a}T\big)\simeq 1$ holds true, to the right it is no longer valid.
@@ III.1 Likelihood analysis
where the observational data $\mathbf{d}=\mathbf{s}+\mathbf{n}$ consists of the axion signal $\mathbf{s}$ and noise $\mathbf{n}$. $\mathbf{\widetilde{d}}$ denotes $\mathbf{d}$ with the intrinsic effect subtracted, $\mathbf{\Theta}=\{g,m_{a}\}$ denotes the set of model parameters and $\mathbf{\Sigma}=\mathbf{\Sigma}^{(s)}+\mathbf{\Sigma}^{(n)}$ is the cross-correlation matrix. In practice, each satellite in the network typically produces on the order of $10^{4}\sim 10^{5}$ signals. As a result, $\mathbf{\Sigma}$ in Eq. (III.3) becomes prohibitively large, and the associated computational burden severely degrades the efficiency of solving the system. Fortunately, [33, 14, 19] demonstrate that $g_{95\%}$ can be determined from the condition: $\big\langle q\big\rangle=2.71$, where
$\big\langle q\big\rangle$ $=\frac{1}{2}\sum_{i,j}\frac{1}{\lambda_{i}\lambda_{j}}\mathrm{Tr}\left({\textstyle\sum_{ij}^{(s)}}{\textstyle\sum_{ji}^{(s)}}\right),$ (III.4)
$\mathrm{Tr}\left({\textstyle\sum_{ii}^{(s)}}{\textstyle\sum_{ii}^{(s)}}\right)$ $=4\frac{g_{a\gamma}^{4}}{m_{a}^{4}}N_{i}^{2}\rho_{\text{DM}}^{2}\Big[1-\mathrm{cos}\big(m_{a}L_{i}\big)\mathrm{sinc}\big(y_{ei}\big)\Big]^{2},$
$\mathrm{Tr}\left({\textstyle\sum_{ij}^{(s)}}{\textstyle\sum_{ij}^{(s)}}\right)$ $=\frac{g_{a\gamma}^{4}}{m_{a}^{4}}N_{i}N_{j}\rho_{\text{DM}}^{2}\Big[1+\mathrm{sinc}^{2}\big(y_{ij}\big)$
$\quad+2\mathrm{cos}\big(m_{a}\big(L_{i}-L_{j}\big)\big)\mathrm{sinc}\big(y_{ij}\big)$
$\quad+f\big(y_{ei},y_{ej}\big)\Big].$
The factor $f\big(y_{ei},y_{ej}\big)$ in the above expression is given by
$f\big(y_{ei},y_{ej}\big)$ $=-2\Big[\mathrm{cos}\big(m_{a}L_{i}\big)\big(\mathrm{sinc}\big(y_{ei}\big)+\mathrm{sinc}\big(y_{ej}\big)\mathrm{sinc}\big(y_{ij}\big)\big)$ (III.5)
$\quad+\mathrm{cos}\big(m_{a}L_{j}\big)\big(\mathrm{sinc}\big(y_{ej}\big)+\mathrm{sinc}\big(y_{ei}\big)\mathrm{sinc}\big(y_{ij}\big)\big)\Big]$
@@
$\quad+2\mathrm{cos}\big(m_{a}\big(L_{i}+L_{j}\big)\big)\mathrm{sinc}\big(y_{ei}\big)\mathrm{sinc}\big(y_{ej}\big)\Big].$
@@
Here $L_{i}$ is the distance between the $i$-th transmitting satellite and the receiving satellite, $y_{ij}\equiv\left|\mathbf{r}_{i}-\mathbf{r}_{j}\right|/\lambda_{\text{de}}$ and $y_{ei}\equiv\left|\mathbf{r}_{i}\right|/\lambda_{\text{de}}$ are the dimensionless distance parameters and $N_{i}$ denotes the total number of pulse signals emitted by the $i$-th transmitting satellite within the coherence time $\tau_{c}$. $\mathrm{sinc}\big(y\big)\equiv\mathrm{sin}\big(y\big)/y$ is the sinc function. [33] notes that the factor $f\big(y_{ei},y_{ej}\big)$ is negligible when $y_{ei}\gg 1$, but must be retained when $y_{ei}\leq 1$. For pulsars located at cosmological distances, the condition $y_{ei}\gg 1$ is typically satisfied; in contrast, for the APPA considered in this work, the satellite network occupies a single coherent axion domain, the parameter $y_{ei}\leq 1$ and in practice $y_{ei}\ll 1$, thus $\mathrm{sinc}\big(y_{ei}\big)\simeq 1$. Consequently, $f\big(y_{ei},y_{ej}\big)$ plays an important role in the analysis and cannot be neglected.
@@
In the numerical simulation, we assume a total of $N_{\mathrm{sat}}=6$ transmitting satellites. The receiving satellite is placed at the center, while the transmitting satellites are randomly distributed in all directions with their distances $L_{i}$. White noise is generated according to Eq. (III.1) and Eq. (III.2), with $\sigma_{\text{err}}=0.3^{\circ}$. For each transmitting satellite ($i=1,2,\dots,6$), the number of pulse signals is set to $N_{i}=10^{5}$. Here, using Eq. (III.4) we show the impact of the axion mass $m_{a}$ and the characteristic scale $L$ (the average distance from the transmitting satellite to the receiving satellite) of the satellite network on the 95% C.L. upper limits $g_{95\%}$, as illustrated in Fig. 3 and Fig. 4.
@@
Figure 3: 95% C.L. upper limit $g_{95\%}$ on the function $g_{a\gamma}\big(m_{a}\big)$, which is calculated from $\big\langle q\big\rangle=2.71$ and Eq. (III.4). Color coding: Red, $L=2.5\times 10^{6}\mathrm{km}$ (LISA [3]); Magenta, $L=2.6\times 10^{8}\mathrm{km}$ (ASTROD-GW [37]); Pink, $L=7.5\times 10^{8}\mathrm{km}$ (Jupiter’s orbit, 5AU).
@@
Figure 4: 95% C.L. upper limit $g_{95\%}$ on the function $g_{95\%}\big(L\big)$, which is calculated from $\big\langle q\big\rangle=2.71$ and Eq. (III.4). The transition point at which $g_{95\%}$ changes from being sensitive to insensitive to the distance is approximately given by the corresponding Compton wavelength $\lambda_{\text{com}}$ of axion. Color coding: Cyan, $m_{a}=10^{-14}\mathrm{eV}$; Green, $m_{a}=10^{-16}\mathrm{eV}$; Blue, $m_{a}=10^{-18}\mathrm{eV}$.
@@
Fig. 3 indicates that a larger spatial scale of the satellite network is more favorable for detecting axion signals, providing better sensitivity to light axions. Unlike observations using natural pulsars, in our work Eq. (III.4) incorporates the effect of $f$, for light axions with Compton wavelengths $\lambda_{\text{com}}\geq L$, $g_{95\%}$ shows no significant variation on the order of magnitude. A similar behavior can also be seen in Fig. 4, which shows the 95% C.L. upper limits $g_{95\%}$ on the function $g_{a\gamma}\big(L\big)$ for different $m_{a}$. In this figure, the characteristic scale $L$ of the satellite network ranges from $10^{5}\mathrm{km}$ to $10^{9}\mathrm{km}$, covering scales from the Earth–Moon distance to the Jupiter–Sun distance. We find that for lighter axions, the 95% C.L. upper limit is more sensitive to $L$. A larger $L$ leads to a stronger constraint on $g_{a\gamma}$. Likewise, a saturation scale exists: once $L$ exceeds the Compton wavelength $\lambda_{\text{com}}$ of axion, further increasing does not improve the 95% C.L. upper limit.

### v3

@@ Abstract
The approaches to searching for axion-like signals based on pulsars include observations with pulsar timing arrays (PTAs) and pulsar polarization arrays (PPAs). However, these methods are limited by observational uncertainties arising from multiple unknown and periodic physical effects, which substantially complicate subsequent data analysis. To mitigate these issues and improve data fidelity, we propose the Artificial Pulsar Polarization Arrays (APPA): a satellite network comprising multiple pulsed signal transmitters and a dedicated receiver satellite. To constrain the axion-photon coupling parameter $g_{a\gamma}$, we generate simulated observations using Monte Carlo methods and investigate the sensitivity of APPA using two complementary approaches: Likelihood analysis and frequentist analysis. Simulations indicate that for the axion mass range of $10^{-22}-10^{-18}$ eV, APPA yields a tighter upper limit on $g_{a\gamma}$ (at the 95% C.L.) than conventional ground-based observations, while also achieving superior detection sensitivity. Moreover, a larger spatial distribution scale of the satellite network corresponds to a greater advantage in detecting axions with lighter masses.
@@ II Axion induced birefringence effect
In [35], the authors considered the occupation number of a scalar axion field and estimated it under standard local halo assumptions. Assuming a characteristic dark-matter velocity $v\sim 10^{-3}c$, and a local dark matter energy density $\rho_{\text{DM}}=0.4$GeV/cm3 in the vicinity of the Solar System, the occupation number can be estimated by
@@
Meanwhile, with the typical velocity $v\sim 10^{-3}c$, we compile a list of useful physical quantities, including the Compton frequency $f_{\mathrm{c}}$, the de Broglie wavelength $\lambda_{\mathrm{de}}$ and the coherence time $\tau_{\mathrm{coh}}$,
[…]
For a given axion mass, if the condition $\tau_{\mathrm{coh}}\gg T_{\text{obs}}$ is satisfied, then the observed axion field amplitude behaves stochastically.
As shown in [22, 15, 14], within a single coherence domain the amplitude distribution of the axion field follows a Rayleigh distribution,
$a_{0}=\frac{\sqrt{2\rho_{\text{DM}}}}{m_{a}}\alpha,\qquad p\big(\alpha\big)\sim\alpha e^{-\frac{\alpha^{2}}{2}}.$ (II.9)
By combining Eq. (II.4), Eq. (II.7) and Eq. (II.9), we obtain the expression for the general case
$\Delta\phi=\phi_{a}\mathrm{cos}\big(m_{a}t+\varphi_{a}\big),$ (II.10)
@@
$\phi_{a}$ $=\frac{g_{a\gamma}}{\sqrt{2}m_{a}}\left(\rho_{o}\alpha_{o}^{2}+\rho_{s}\alpha_{s}^{2}-2\sqrt{\rho_{o}\rho_{s}}\alpha_{o}\alpha_{s}\mathrm{cos}\theta_{T}\right)^{\frac{1}{2}},$ (II.11)
$\varphi_{a}$ $=\arctan\frac{\sqrt{\rho_{o}}\alpha_{o}\sin\varphi_{o}+\sqrt{\rho_{s}}\alpha_{s}\sin\big(m_{a}T-\varphi_{s}\big)}{\sqrt{\rho_{o}}\alpha_{o}\cos\varphi_{o}-\sqrt{\rho_{s}}\alpha_{s}\cos\big(m_{a}T-\varphi_{s}\big)},$
with $\theta_{T}\equiv m_{a}T+\varphi_{o}-\varphi_{s}$, and $T$ is the signal propagation time.
@@ III Methods of analysis
2. Uncertainty in local dark matter energy density. As indicated in Eq. (II.11), the measurement of the axion–photon coupling parameter $g_{a\gamma}$ requires knowledge of the dark matter energy densities $\rho_{s}$ near the pulsar and $\rho_{o}$ near the Solar System. However, $\rho_{s}$ is difficult to determine observationally and is subject to large uncertainties, which ultimately limit the precision of the derived constraints.
3. Randomness of phase. For pulsars, the pulsar and the observer often reside in different axion coherent regions. This means the phase $\varphi_{s}$ of the axion field at a pulsar is inherently random and indeterminate, which prevents the signal propagation time $T$ from being used to reliably determine the value of $\theta_{T}$ in Eq. (II.11). Meanwhile, the phase of the pulsar signals cannot be well determined, introducing uncertainty into practical research and analysis.
@@
To address these observational challenges at their source, we extend the concept proposed in [2] by introducing the APPA, a network of space-based satellites equipped with ultra precise clocks and capable of regularly transmitting pulsed signals to a dedicated receiving satellite. The transmitted signals are strictly periodic, with a timing uncertainty of $\sigma_{t}=\xi_{\delta t}\delta t$ [2], and carry preconfigured polarization information. As illustrated in Fig. 1, the receiving satellite, positioned at the center of the transmitting network, continuously collects APPA signals and can accurately measure both the distance to and the direction of each transmitting satellite. We assume that the entire satellite network is deployed within the solar system, therefore it can be treated as residing within a single coherent axion field, implying $\varphi_{o}=\varphi_{s}=\varphi$, $\rho_{o}=\rho_{s}=\rho_{\text{DM}}$ and $\alpha_{o}=\alpha_{s}=\alpha$ in Eq. (II.11). Consequently, a significant fraction of environmental disturbances and intrinsic signal uncertainties are effectively eliminated, and Eq. (II.11) reduces to
$\phi_{a}$ $=\frac{g_{a\gamma}}{m_{a}}\sqrt{\frac{\rho_{\text{DM}}}{2}}\alpha\big[2\left(1-\mathrm{cos}\left(m_{a}T\right)\right)\big]^{\frac{1}{2}},$ (III.1)
$\varphi_{a}$ $=\arctan\frac{\sin\varphi+\sin\big(m_{a}T-\varphi\big)}{\cos\varphi-\cos\big(m_{a}T-\varphi\big)}.$
@@
$\lambda=\big(EF\sigma_{\text{err}}\big)^{2}+EQ^{2}.$ (III.2)
Here, $\sigma_{\text{err}}$ denotes the instrumental measurement error, which we take to be $\sigma_{\text{err}}=0.3^{\circ}$. Eq. (III.2) accounts for measurement uncertainties, which are modeled by two parameters $EF$ and $EQ$, with $EF$ being the white noise scaling factor and $EQ$ being the additional white noise. Both of them are free parameters. For more details on white noise, please refer to [61, 47, 24, 7, 60, 46]. Here, the prior distributions are specified as follows:
$P\big(EF\big)$ $=\mathcal{N}\big[1,1\big],$ (III.3)
@@
First, we show the main conclusions of this work in advance. We present a comparison between the simulation results of this work and existing observational constraints in a single figure (see Fig. 2). Owing to the long observational baseline $T_{\text{obs}}$ and the highly stable signal emission cadence $\Delta t$, APPA is theoretically capable of probing an exceptionally broad range of axion induced oscillation frequencies. However, as shown in Fig. 2, although APPA theoretically possesses an extremely wide detection range, the domain where it truly demonstrates an advantage in axion signal detection is limited to $m_{a}\sim\mathcal{O}\big(10^{-22}–10^{-18}\big)$eV, both the 95% C.L. upper limit $g_{95\%}$ derived from likelihood analysis and the detection sensitivity $g_{a\gamma}$ obtained via frequentist analysis highlight the significant advantages of the APPA framework over current ground-based observations.
Figure 2: Comparison between the APPA simulation results and existing observational constraints. Color coding: Red, CAST [5]; Magenta, PPA [61]; Orange, EPTA [46]; Cyan, PSR J0437–4715 [13]; Green, PPTA [14]; Blue, likelihood analysis (L=5L=5AU), we shows that APPA has a detection advantage within $m_{a}\sim\mathcal{O}\big(10^{-22}–10^{-18}\big)$eV and exhibits comparable sensitivity to axions with different masses; Purple, frequentist analysis, the detection sensitivity is set to $\phi_{a}=0.003^{\circ}$ in Eq. (III.13), consistent with current observational limits on linear polarization angle rotation, with minor fluctuations attributed to instrumental noise. The black dashed line marks the critical boundary, to the left of this boundary, the condition $\mathrm{cos}\big(m_{a}T\big)\simeq 1$ holds true. To the right it is no longer valid, we adopt $m_{a}T=0.1$ as the reference criterion.
@@ III.1 Likelihood analysis
where the observational data $\mathbf{d}=\mathbf{s}+\mathbf{n}$ consists of the axion signal $\mathbf{s}$ and noise $\mathbf{n}$, $\mathbf{\widetilde{d}}$ denotes $\mathbf{d}$ with the intrinsic effect subtracted, $\mathbf{\Theta}=\{g_{a\gamma},m_{a}\}$ denotes the set of model parameters and $\mathbf{\Sigma}=\mathbf{\Sigma}^{(s)}+\mathbf{\Sigma}^{(n)}$ is the $N_{i}\times N_{j}$ cross-correlation matrix. In practice, each satellite in the network typically generates $N_{i}$ on the order of $10^{4}\sim 10^{5}$ signals. According to [36, 16, 21], if we given a mass $m_{a}$, the exclusion limit for $g_{a\gamma}$ is set by a test statistic,
$q\big(g_{a\gamma},m_{a}\big)\equiv 2\big[\mathrm{ln}\mathcal{L}\big(\hat{g}_{a\gamma},m_{a}|\mathbf{d}\big)-\mathrm{ln}\mathcal{L}\big(g_{a\gamma},m_{a}|\mathbf{d}\big)\big],$ (III.5)
where $\hat{g}_{a\gamma}$ maximizes $\mathcal{L}\big(g_{a\gamma},m_{a}|\mathbf{d}\big)$. With $q(g_{a\gamma},m_{a})=0$ for $g_{a\gamma}<\hat{g}_{a\gamma}$, the upper limit at 95% C.L. $g_{95\%}$ is given by $q\big(g_{95\%},m_{a}\big)=2.71$.
As a result, $g_{95\%}$ can be determined from the condition: $\big\langle q\big\rangle=2.71$ [36], where
$\big\langle q\big\rangle=\frac{1}{2}\sum_{i,j}\frac{1}{\lambda_{i}\lambda_{j}}\mathrm{Tr}\left(\mathbf{\Sigma}^{(s)}_{ij}\mathbf{\Sigma}^{(s)}_{ji}\right)$. (III.6)
When $i=j$, the trace term in Eq. (III.6) is
$\mathrm{Tr}\left(\mathbf{\Sigma}^{(s)}_{ii}\mathbf{\Sigma}^{(s)}_{ii}\right)$ $=4\frac{g_{a\gamma}^{4}}{m_{a}^{4}}N_{i}^{2}\rho_{\text{DM}}^{2}\Big[1-\mathrm{cos}\big(m_{a}T_{i}\big)\mathrm{sinc}\big(y_{ei}\big)\Big]^{2},$ (III.7)
where $\mathbf{\Sigma}^{(s)}_{ii}$ represents the auto-correlation of the $i$-th source. When $i\neq j$, the trace term in Eq. (III.6) is
$\mathrm{Tr}\left(\mathbf{\Sigma}^{(s)}_{ij}\mathbf{\Sigma}^{(s)}_{ji}\right)$ $=\frac{g_{a\gamma}^{4}}{m_{a}^{4}}N_{i}N_{j}\rho_{\text{DM}}^{2}\Big[1+\mathrm{sinc}^{2}\big(y_{ij}\big)$ (III.8)
$\quad+2\mathrm{cos}\big(m_{a}\big(T_{i}-T_{j}\big)\big)\mathrm{sinc}\big(y_{ij}\big)$
$\quad+f\big(y_{ei},y_{ej}\big)\Big],$
where $\mathbf{\Sigma}^{(s)}_{ij}$ represents the cross-correlation between the $i$-th and $j$-th sources. The factor $f\big(y_{ei},y_{ej}\big)$ in the above expression is given by
$f\big(y_{ei},y_{ej}\big)$ $=-2\Big[\mathrm{cos}\big(m_{a}T_{i}\big)\big(\mathrm{sinc}\big(y_{ei}\big)+\mathrm{sinc}\big(y_{ej}\big)\mathrm{sinc}\big(y_{ij}\big)\big)$ (III.9)
$\quad+\mathrm{cos}\big(m_{a}T_{j}\big)\big(\mathrm{sinc}\big(y_{ej}\big)+\mathrm{sinc}\big(y_{ei}\big)\mathrm{sinc}\big(y_{ij}\big)\big)\Big]$
@@
$\quad+2\mathrm{cos}\big(m_{a}\big(T_{i}+T_{j}\big)\big)\mathrm{sinc}\big(y_{ei}\big)\mathrm{sinc}\big(y_{ej}\big)\Big].$
@@
Here $T_{i}=L_{i}/c$ is the signal travel time from the $i$-th transmitting satellite to the receiving satellite, and $L_{i}$ is the distance, respectively. $y_{ij}\equiv\left|\mathbf{r}_{i}-\mathbf{r}_{j}\right|/\lambda_{\text{de}}$ and $y_{ei}\equiv\left|\mathbf{r}_{i}\right|/\lambda_{\text{de}}$ are the dimensionless distance parameters and $N_{i}$ denotes the total number of pulse signals emitted by the $i$-th transmitting satellite within the coherence time $\tau_{\mathrm{coh}}$. $\mathrm{sinc}\big(y\big)\equiv\mathrm{sin}\big(y\big)/y$ is the sinc function. In [36], it has been noted that the factor $f\big(y_{ei},y_{ej}\big)$ is negligible when $y_{ei}\gg 1$, but must be retained when $y_{ei}\leq 1$. For pulsars located at cosmological distances, the condition $y_{ei}\gg 1$ is typically satisfied; in contrast, for the APPA considered in this work, the satellite network occupies a single coherent axion domain, the parameter $y_{ei}\leq 1$ and in practice $y_{ei}\ll 1$, thus $\mathrm{sinc}\big(y_{ei}\big)\simeq 1$. Consequently, $f\big(y_{ei},y_{ej}\big)$ plays an important role in the analysis and cannot be neglected.
@@
In the numerical simulation, we assume a total of $N_{\mathrm{sat}}=6$ transmitting satellites. The receiving satellite is placed at the center, while the transmitting satellites are randomly distributed in all directions with their distances $L_{i}$. White noise is generated according to Eq. (III.2) and Eq. (III.3), with $\sigma_{\text{err}}=0.3^{\circ}$. For each transmitting satellite ($i=1,2,\dots,6$), the number of pulse signals is set to $N_{i}=10^{5}$. Here, using Eq. (III.6) we show the impact of the axion mass $m_{a}$ and the characteristic scale $L$ (the average distance from the transmitting satellite to the receiving satellite) of the satellite network on the 95% C.L. upper limits $g_{95\%}$, as illustrated in Fig. 3 and Fig. 4.
@@
Figure 3: 95% C.L. upper limit $g_{95\%}$ on the function $g_{a\gamma}\big(m_{a}\big)$, which is calculated from $\big\langle q\big\rangle=2.71$ and Eq. (III.6). Color coding: Red, $L=2.5\times 10^{6}\mathrm{km}$ (LISA [3]); Magenta, $L=2.6\times 10^{8}\mathrm{km}$ (ASTROD-GW [40]); Pink, $L=7.5\times 10^{8}\mathrm{km}$ (Jupiter’s orbit, 5AU).
@@
Figure 4: 95% C.L. upper limit $g_{95\%}$ on the function $g_{95\%}\big(L\big)$, which is calculated from $\big\langle q\big\rangle=2.71$ and Eq. (III.6). The transition point at which $g_{95\%}$ changes from being sensitive to insensitive to the distance is approximately given by the corresponding Compton wavelength $\lambda_{\text{c}}$ of axion. Color coding: Cyan, $m_{a}=10^{-15}\mathrm{eV}$; Green, $m_{a}=10^{-16}\mathrm{eV}$; Blue, $m_{a}=10^{-18}\mathrm{eV}$.
@@
Fig. 3 indicates that a larger spatial scale of the satellite network is more favorable for detecting axion signals, providing better sensitivity to light axions. Unlike observations using natural pulsars, in our work Eq. (III.6) incorporates the effect of $f$, for light axions with Compton wavelengths $\lambda_{\text{c}}\geq L$, $g_{95\%}$ shows no significant variation on the order of magnitude. A similar behavior can also be seen in Fig. 4, which shows the 95% C.L. upper limits $g_{95\%}$ on the function $g_{a\gamma}\big(L\big)$ for different $m_{a}$. In this figure, the characteristic scale $L$ of the satellite network ranges from $10^{5}\mathrm{km}$ to $10^{9}\mathrm{km}$, covering scales from the Earth–Moon distance to the Jupiter–Sun distance. We find that for lighter axions, the 95% C.L. upper limit is more sensitive to $L$. A larger $L$ leads to a stronger constraint on $g_{a\gamma}$. Likewise, a saturation scale exists: once $L$ exceeds the Compton wavelength $\lambda_{\text{c}}$ of axion, further increasing does not improve the 95% C.L. upper limit.